3.7 An insurance company is looking to replace its legacy claims payment system. The Technology Steering Committee is willing to explore three alternative systems. Table EX 3.7 summarizes the cost savings for each system under potential peak claims volumes. Table EX 3.7 Payoff: Cost Savings (in $000s) for Peak ClaIMS: see table below a. The most recent system implementation was riddled with data conversion issues and end-user resistance. Accordingly, members of the Technology Steering Committee are relatively pessimistic decision makers. Which system would be selected under a Hurwitz value of 0.2? b. Which system would be selected under minimax regret? c. Which system would be selected using the Laplace strategy solution? d. The company's information technology (IT) department submitted a proposal to develop an in-house solution for claims payment. This in-house solution is expected to provide cost savings of $25,000 for five million claims, $75,000 for ten million claims, and $120,000 for fifteen million claims. Determine if the system selection changes in parts (a) through (c) when the in-house system is added as an alternative.

Payoff: Cost Savings (in $000s) for Peak Claims Volumes System

System 5,000,000 CLAIMS 10,000,000 CLAIMS 15,000,000 CLAIMS A

A 35 70 105

B (50) 85 170

C 10 25 75
Answer #1

a) b) c)

A 35 70 2 System 3 A 4 B 5 C 5,000,000 Claims 10,000,000 Claims 15,000,000 Claims Min payoff Hurwicz payoff Laplace payoff 10

Formula

5,000,000 Claims 10,000,000 Claims 2 System 3 A 4 B 5 C 15,000,000 Claims 105 170 Min payoff =MIN(B3:03) =MIN(B4:04) =MIN(B5:

d)

With the in-house option,

F G H -50 Laplace payoff 49 -6 68.33333333 23 36.66666667 44 73.33333333 85 2 System 5,000,000 Claims 10,000,000 Claims 15,00

Formula

ДА 5,000,000 Claims 10,000,000 Claims 15,000,000 Claims 105 85 170 -50 10 Min payoff =MIN(B3:D3) =MIN(B4:04) =MIN(B5:05) =MIN

As seen. decision changes to In-house for Laplace and Minimax regret criteria